# TruthTensor: Evaluating LLMs Human Imitation through Prediction Market Drift and Holistic Reasoning

### v1

@@ Abstract
Evaluating language models and AI agents remains fundamentally challenging because static benchmarks fail to capture real-world uncertainty, distribution shift, and the gap between isolated task accuracy and human-aligned decision-making under evolving conditions. This paper introduces TruthTensor, a novel, reproducible evaluation paradigm that measures Large Language Models (LLMs) not only as prediction engines but as human-imitation systems operating in socially-grounded, high-entropy environments. Building on forward-looking, contamination-free tasks, our framework anchors evaluation to live prediction markets and combines probabilistic scoring to provide a holistic view of model behavior. TruthTensor complements traditional correctness metrics with drift-centric diagnostics and explicit robustness checks for reproducibility. It specify human vs. automated evaluation roles, annotation protocols, and statistical testing procedures to ensure interpretability and replicability of results. In experiments across 500+ real markets (political, economic, cultural, technological), TruthTensor demonstrates that models with similar forecast accuracy can diverge markedly in calibration, drift, and risk-sensitivity, underscoring the need to evaluate models along multiple axes (accuracy, calibration, narrative stability, cost, and resource efficiency). TruthTensor therefore operationalizes modern evaluation best practices, clear hypothesis framing, careful metric selection, transparent compute/cost reporting, human-in-the-loop validation, and open, versioned evaluation contracts, to produce defensible assessments of LLMs in real-world decision contexts. We publicly release TruthTensor at https://truthtensor.com.
@@ 1.2. Key Differentiators
  • Token Constraint Awareness: The framework accounts for token budget limitations, measuring how reasoning quality degrades under resource constraints.
@@ 2.5. Drift: The Central Evaluation Dimension
  • Narrative Consistency Score: Measures the coherence of reasoning traces across time points, penalizing contradictory justifications.
@@ 3.1. Instruction Locking and Prompt Specification Layer
  • a forecasting target (e.g., “Will candidate X win the upcoming election?”),
  • a time horizon and sampling cadence,
@@ 3.2. Baseline Construction Layer
TruthTensor constructs baseline models that provide reference points independent of rolling-window calibration or model-specific training histories. Baselines include:
  • Market Baseline: Uses current market-implied probabilities as forecasts, representing aggregated human expectations.
  • Uniform Baseline: Assigns equal probability to all outcomes, representing a null model with no information.
  • Historical Baseline: Uses historical event frequencies within categories, representing base-rate reasoning.
  • Simple LLM Baselines: Smaller, simpler models (GPT-2 scale) that provide reference points for comparison.
These baselines are not based on rolling-window calibration from model outputs. Instead, they provide independent reference points that enable fair comparison across models with different architectures, training data, and calibration histories. This design choice ensures that evaluation focuses on model capabilities rather than calibration method differences.
@@ 4.3. Token Constraint Evaluation
  • how reasoning trace quality changes with token limits,
@@ 5.8. Behavioral and Temporal Diagnostics
Figure 2 compares model-driven outcomes to a historical market baseline (left) and to realized live trading outcomes (right). The adjusted P&L curve isolates model signal from market drift by subtracting baseline movement, while the realized P&L reflects execution under real trading constraints. The divergence between these curves reveals that some models generate strong directional information (high adjusted P&L) but lose value through volatility, slippage, or mis-timed adjustments in live conditions. Models that maintain flatter realized curves exhibit more stable temporal belief updates, even when their raw signal is weaker.
@@
Figure 6 measures how often models produce decisions that improve predictive alignment. Claude-Sonnet-4.5, Grok-4, and DeepSeek-Chat-v3.1 achieve the highest success counts, indicating reliable local updates. Kimi-K2-Thinking, despite its long reasoning chains, exhibits the lowest success frequency, showing that verbosity does not necessarily translate into effective probabilistic correction.
Together these diagnostics reveal that model quality in live forecasting is governed not only by final accuracy but by how beliefs evolve. High-capacity models generate stronger raw signals but also exhibit greater volatility and adjustment depth. More compact models trade expressive power for temporal coherence, yielding steadier probability trajectories. TruthTensor makes these behavioral differences measurable, allowing models to be compared not just by what they predict, but by how they reason over time.
@@ 6. Conclusion
This work presents TruthTensor as a new class of benchmark for evaluating large language models in forward-looking decision environments. By embedding LLM agents directly into live prediction markets, the framework observes not only what models predict, but how they revise their beliefs over time under real uncertainty. Across nearly one million probability updates and eight simultaneously deployed frontier-scale models, we show that performance cannot be summarized by a single score. Models that achieve similar predictive accuracy often differ substantially in calibration, temporal stability, and alignment with market information, revealing qualitatively different reasoning dynamics.

### v2

# TruthTensor: Evaluating LLMs through Human Imitation on Prediction Market under Drift and Holistic Reasoning

@@ Abstract
Evaluating language models and AI agents remains fundamentally challenging because static benchmarks fail to capture real-world uncertainty, distribution shift, and the gap between isolated task accuracy and human-aligned decision-making under evolving conditions. This paper introduces TruthTensor, a novel, reproducible evaluation paradigm that measures reasoning models not only as prediction engines but as human-imitation systems operating in socially-grounded, high-entropy environments. Building on forward-looking, contamination-free tasks, our framework anchors evaluation to live prediction markets and combines probabilistic scoring to provide a holistic view of model behavior. TruthTensor complements traditional correctness metrics with drift-centric diagnostics and explicit robustness checks for reproducibility. It specify human vs. automated evaluation roles, annotation protocols, and statistical testing procedures to ensure interpretability and replicability of results. In experiments across 500+ real markets (political, economic, cultural, technological), TruthTensor demonstrates that models with similar forecast accuracy can diverge markedly in calibration, drift, and risk-sensitivity, underscoring the need to evaluate models along multiple axes (accuracy, calibration, narrative stability, cost, and resource efficiency). TruthTensor therefore operationalizes modern evaluation best practices, clear hypothesis framing, careful metric selection, transparent compute/cost reporting, human-in-the-loop validation, and open, versioned evaluation contracts, to produce defensible assessments of LLMs in real-world decision contexts. We publicly release TruthTensor at https://truthtensor.com.
@@ 3.1. Instruction Locking and Prompt Specification Layer
  • a binary forecasting target which is either a yes or a no,
  • a time horizon that reflects cycling and tuning (e.g. 466 cycle representing api call for each model),
@@ 3.2. Baseline Construction Layer
In this study, the baseline is defined exclusively as the market baseline, constructed directly from live prediction market prices. All baseline information used throughout the paper is derived solely from the market signals summarized in Table 2, which therefore functions as the definitive reference frame for evaluation. Each row corresponds to a model or agent evaluated during the same market window, while the columns capture both market-aligned behavioral signals and agent-level outcomes relative to that market.
  • P&L (profit and loss) quantify the economic consequence of these deviations when agent decisions are hypothetically executed against the market. Positive P&L values indicate that an agent’s probability estimates improve upon the market consensus in expectation, while negative values reflect miscalibration or delayed information assimilation.
  • Unique Users and Agent Count provide context on interaction scale and deployment intensity during the evaluation window. Higher user counts indicate broader exposure to market conditions, while agent count reflects the degree of parallel decision-making evaluated under the same baseline.
  • Average Input Tokens and Average Output Tokens describe computational characteristics of each agent. These values contextualize efficiency trade-offs: for instance, agents with comparable P&L but significantly lower token usage demonstrate more efficient reasoning relative to the same market baseline.
@@ 5.8. Behavioral and Temporal Diagnostics
Figure 2 illustrates both the adjusted cumulative performance of the model relative to historical baselines (left) and the realized cumulative trading profit and loss over time during the TruthTensor experiment (right). The left panel demonstrates that the proposed agent consistently outperforms market, uniform, and historical benchmark strategies, exhibiting a monotonic upward trajectory that reflects stable probabilistic calibration and sustained positive edge extraction across evaluation rounds. In contrast, the right panel captures realized market exposure under the fixed 30-position portfolio constraint, highlighting periods of volatility in which some strategies incur drawdowns due to miscalibration, drift amplification, or high-risk event exposure. The divergence between the adjusted and realized curves emphasizes the importance of separating intrinsic forecasting quality from market-execution noise. Notably, the dominant model maintains controlled losses during turbulent intervals and recovers steadily, whereas weaker baselines exhibit prolonged negative P&L, validating the robustness of the drift-aware, calibration-adjusted, and risk-constrained decision framework.
@@
Figure 6 measures the frequency with which models successfully return valid system-prompt outputs within predefined decision-time and latency constraints. Claude-Sonnet-4.5, Grok-4, and DeepSeek-Chat-v3.1 achieve the highest success counts, indicating reliable local updates. Kimi-K2-Thinking, despite its long reasoning chains, exhibits the lowest success frequency, showing that verbosity does not necessarily translate into effective probabilistic correction. These diagnostics reveal that model quality in live forecasting is governed not only by final accuracy but by how beliefs evolve. High-capacity models generate stronger raw signals but also exhibit greater volatility and adjustment depth. More compact models trade expressive power for temporal coherence, yielding steadier probability trajectories. TruthTensor makes these behavioral differences measurable, allowing models to be compared not just by what they predict, but by how they reason over time.
@@ 6. Conclusion
This work presents TruthTensor as a new class of benchmark for evaluating large language models in forward-looking decision environments. By embedding LLM agents directly into live prediction markets, the framework observes not only what models predict, but how they revise their beliefs over time under real uncertainty. Across nearly one million probability updates and eight simultaneously deployed frontier-scale models, we show that performance cannot be summarized by a single score.
@@
Recent work [22] demonstrates that LLM agents can be used as digital twins to simulate human behaviour and evaluate agentic systems by comparing agent-generated interactions with real human responses. Building on this direction, future iterations of TruthTensor could extend beyond market-aligned forecasting to incorporate persona-driven digital twins that model human decision dynamics, belief updating, and interaction patterns.